A 2-C charge and a 4-C charge attract each other with 10 N of force. How much will a 2-C charge and a 12-C charge attract each other when placed the same distance apart?
1
Expert's answer
2018-03-07T11:40:07-0500
The Coulomb’s law states
F_12=k (q_1 q_2)/r^2
F_13=k (q_1 q_3)/r^2
Since
q_3=3q_2
So
F_13=3F_12=3×10=30 N